show that the two vectors a = 9i + j - 4k, b = 3i - 7j + 5k are mutually perpendicular
Answer by tommyt3rd(5050)

Vectors are perpendicular if their dot product is 0:
(9)(3)+(1)(-7)+(-4)(5)=27-7-20=0



:)